1) A car initially moves at 60 mph and then slows down to 45 mph after traveling a distance of 40 m. Determine its acceleration, assuming it is constant, in units of m/s^2.


a) 20 m/s^2
b) 5.00 m/s^2
c) 15 m/s^2
d) -3.94 m/s^2
Answer #1

Initially car moves at u= 60mph = 1609.35×60/3600= 26.82m/s

Same Final speed of car v= 45mph= 20.11m/s

Total distance travelled d= 40m.

Let acceleration is a

Using formula

v² - u² = 2ad

(20.11)² - (26.82)² = 2×a×40

404.41 - 719.31 = 2×a×40

So a = 314.9/80

So acceleration of car is -3.94m/s²